Zari:

What is the length of the hypotenuse of a right triangle if each of the two legs is 7 units Square root of 14 7 Square root of 98 49

Vocaloid:

If you have two equal legs you can simply multiply length of one leg * sqrt(2) to get the hypotenuse In your case, 7*sqrt(2) = ?

Zari:

is it 14 or 7 ?

Vocaloid:

No not quite Remember, the 2 is inside the square root We can rewrite 7 as sqrt(49) since the square root of 49 is 7 So now we have sqrt(49) * sqrt(2) = ?

Zari:

98 or 49

Vocaloid:

Remember your square root rules Sqrt(a) * sqrt(b) = sqrt(a*b) So sqrt(49) * sqrt(2) = ?

Zari:

just 49

Vocaloid:

When you have two square roots multiplied together, keep the root and multiply what’s inside the roots So sqrt(49) * sqrt(2) = sqrt(49*2) = sqrt(98)

Zari:

okay